# Quantum Dynamics with Stochastic Non-Hermitian Hamiltonians

### v1

@@ Appendix B Liouvillian spectrum of the Stochastic Dissipative Qubit
The non-trace-preserving state dynamics is then dictated by $\frac{\mathrm{d}}{\mathrm{d}t}|\tilde{\rho})=\tilde{\mathcal{L}}|\tilde{\rho})$ with the vectorized Liouvillian superoperator in (3) given by
~$\begin{split}\tilde{\mathcal{L}}=&-i(\hat{H}_{0}\otimes\hat{\mathbb{1}}-\hat{\mathbb{1}}\otimes\hat{H}_{0}^{\intercal})-(\hat{L}\otimes\hat{\mathbb{1}}+\hat{\mathbb{1}}\otimes\hat{L}^{\intercal})\\ &+\gamma(\hat{L}^{2}\otimes\hat{\mathbb{1}}+\hat{\mathbb{1}}\otimes(\hat{L}^{2})^{\intercal}+2\hat{L}\otimes\hat{L}^{\intercal}).\end{split}$ (20)
Similarly, the Liouvillian for the SDQ (8) is given in matrix form by
@@
with the cubic polynomial $f(\Lambda)=\Lambda^{3}-\Lambda^{2}(A+B)+\Lambda(4+AB)-2B$. To find its roots, we first shift the variable, using $z=\Lambda-(A+B)/3$ and get a depressed cubic, i.e. a cubic equation without quadratic term
@@
Note that we seem to have obtained 6 solutions from a cubic equation. However, three pairs of solutions are the same. To show this, let $V=-\frac{D}{2}\pm\sqrt{\frac{D^{2}}{4}+C^{3}}\equiv U_{m,\pm}^{3}$ denote any nonzero root of the quadratic equation (24). If $V$ is a root then $\frac{-C^{3}}{V}$ is also a root; this implies that if we change +$\rightarrow$-+ → - in $\Lambda_{m,\pm}$ we exchange the terms $U_{m,\pm}$ with the term $-\frac{C}{U_{m,\pm}}$. So it is enough to consider the eigenvalues $\Lambda_{m,+}$.
@@
In the domain of interest, with $\gamma$ and $\Gamma$ positive, $\lambda_{0}$ is the eigenvalue with the maximum real part. This feature is verified for a wide range of parameters in Fig. 4, where the maximum difference of the real part of eigenvalues $\lambda_{\nu}$ for $\nu\geq 1$ and $\lambda_{0}$ is always shown to be negative. The full spectrum can be checked for many different parameters in Fig. 5. This plot allows us to see the large dissipative gap and no oscillatory behavior in the NI phase (a3, a4, b4), as well as the non-zero imaginary components and a small gap in the $\mathcal{PT}$u phase (c1-3), and the intermediate gap and no oscillations in the $\mathcal{PT}$b phase (a1), as well as the transitions between them
Figure 4: Maximum difference of the real part of $\lambda_{\nu},\nu\geq 1$ with $\lambda_{0}$. The difference is always negative for the displayed range of parameters, therefore $\lambda_{0}$ is always the eigenvalue with the largest real part. The transitions between the $\mathcal{PT}$u, $\mathcal{PT}$b, and NI phases are shown in the dashed and solid line as in Fig. 2 of the main text.
@@
$b^{(0)}_{3}=b^{(0)}_{0}\left(1+\lambda_{0}\frac{\lambda_{0}-A}{2}\right)$ (28)
and $b^{(0)}_{1}=-b^{(0)}_{2}=-ib^{(0)}_{0}\lambda_{0}/2$. We choose $b^{(0)}_{0}$ such as to normalize the eigenvector for real $\lambda_{0}$, yielding the eigenvector associated to the stable steady state
@@
Appendix C Bloch sphere dynamics for the Stochastic Dissipative Qubit
@@
Appendix D Comparison of different numerical approaches
We now compare two different numerical resolutions of the SDQ dynamics. The first is to numerically integrate the nonlinear system of differential equations (31)—we do so by a standard 4th-order Runge-Kutta (RK4) method. The second approach is to get the time evolution of the density matrix from the formal solution of the master equation, i.e. $|\tilde{\rho}_{t})=e^{\tilde{\mathcal{L}}t}|\rho_{0})$, where the equation is vectorized to compute the map $e^{\tilde{\mathcal{L}}t}[\bullet]$ simply through matrix exponentiation. Additionally, this method requires normalization of the state $|\rho_{t})=|\tilde{\rho}_{t})/\mathrm{Tr}(\tilde{\rho}_{t})$, where $\mathrm{Tr}(\tilde{\rho}_{t})=(\mathbb{1}|\tilde{\rho}_{t})=\tilde{\rho}_{ee}+\tilde{\rho}_{ff}$ and $\tilde{\rho}_{ee}=\bra{e}\tilde{\rho}_{t}\ket{e}$. To avoid computing a matrix exponential for each time $t$, we Trotterize the evolution as $e^{\tilde{\mathcal{L}}t}=\prod_{n=1}^{N_{\textsc{t}}}e^{\tilde{\mathcal{L}}\Delta_{\textsc{t}}}$, subdividing the evolution in $N_{\textsc{t}}$ steps of length $\Delta_{\textsc{t}}=t_{\rm end}/N_{\textsc{t}}$.
These two approaches are compared in Fig. 7 to compute the purity, showing a perfect agreement. Additionally, the first-order evolution of the purity obtained from $\partial_{t}P$ in (9) also agrees with the simulated evolution.
@@ Appendix D Comparison of different numerical approaches
Appendix E Uhlmann fidelity for a qubit
@@
This expression was used in Fig. 3 to compute the distinguishability between the instantaneous state and the stable steady state.
Appendix F Standard form of the master equation
@@
where $\tilde{\mathcal{L}}$ need not be trace-preserving. Balanced gain and loss is satisfied provided that the normalized density matrix $\hat{\rho}$ evolves according to
@@
However, by relaxing this condition, the general structure of ~$\tilde{\mathcal{L}}[\hat{\rho}]$ is considered. The rate of norm change in this case is
@@
$\frac{d}{dt}\hat{\rho}$ $=$ $-i[\hat{H},\hat{\rho}]+\{\hat{G},\hat{\rho}\}+\sum_{ij=1}^{N^{2}}a_{ij}\hat{F}_{i}\hat{\rho}\hat{F}_{j}^{\dagger}$
$-\Big{(}2{\rm Tr}(\hat{G}\hat{\rho})+\sum_{ij=1}^{N^{2}}a_{ij}{\rm Tr}(\hat{F}_{j}^{\dagger}\hat{F}_{i}\hat{\rho})\Big{)}\hat{\rho}.$
@@
$\frac{d}{dt}\hat{\rho}$ $=$ $-i[\hat{H},\hat{\rho}]+\{\hat{G},\hat{\rho}\}+\sum_{ij=1}^{N^{2}}\gamma_{k}\hat{A}_{k}\hat{\rho}\hat{A}_{k}^{\dagger}$
$-\Big{(}2{\rm Tr}[\hat{G}\hat{\rho}]+\sum_{k=1}^{N^{2}}\gamma_{k}{\rm Tr}[\hat{A}_{k}^{\dagger}\hat{A}_{k}\hat{\rho}]\Big{)}\hat{\rho},$

### v2

@@ Appendix C Liouvillian spectrum of the Stochastic Dissipative Qubit
The non-trace-preserving state dynamics is then dictated by $\frac{\mathrm{d}}{\mathrm{d}t}|\tilde{\rho})=\tilde{\mathcal{L}}|\tilde{\rho})$ with the vectorized Liouvillian superoperator in equation (3) of the main text given by
~$\tilde{\mathcal{L}}=-i(\hat{H}_{0}\otimes\hat{\mathbb{1}}-\hat{\mathbb{1}}\otimes\hat{H}_{0}^{\intercal})-(\hat{L}\otimes\hat{\mathbb{1}}+\hat{\mathbb{1}}\otimes\hat{L}^{\intercal})+\gamma(\hat{L}^{2}\otimes\hat{\mathbb{1}}+\hat{\mathbb{1}}\otimes(\hat{L}^{2})^{\intercal}+2\hat{L}\otimes\hat{L}^{\intercal}).$ (24)
Similarly, the Liouvillian for the SDQ (cf. eq. (8) main text) is given in matrix form by
@@
with the cubic polynomial $f(\Lambda)=\Lambda^{3}-\Lambda^{2}(A+B)+\Lambda(4+AB)-2B$. To find its roots, we first shift the variable, using $z=\Lambda-(A+B)/3$ and get a depressed cubic, i.e. a cubic equation without the quadratic term
@@
Note that we seem to have obtained six solutions from a cubic equation. However, three pairs of solutions are the same. To show this, let $V=-\frac{D}{2}\pm\sqrt{\frac{D^{2}}{4}+C^{3}}\equiv U_{m,\pm}^{3}$ denote any nonzero root of the quadratic equation (28). If $V$ is a root then $\frac{-C^{3}}{V}$ is also a root; this implies that if we change +$\rightarrow$-+ → - in $\Lambda_{m,\pm}$ we exchange the terms $U_{m,\pm}$ with the term $-\frac{C}{U_{m,\pm}}$. So it is enough to consider the eigenvalues $\Lambda_{m,+}$.
@@
Figure 4: Maximum difference of the real part of $\lambda_{\nu},\nu\geq 1$ with $\lambda_{0}$. The difference is always negative for the displayed range of parameters. Therefore, $\lambda_{0}$ is always the eigenvalue with the largest real part. The transitions between the $\mathcal{PT}$u, $\mathcal{PT}$b, and NI phases are shown in the dashed and solid line as in Fig. 2 of the main text.
In the domain of interest, with $\gamma$ and $\Gamma$ positive, $\lambda_{0}$ is the eigenvalue with the maximum real part. This feature is verified for a wide range of parameters in Fig. 4, where the maximum difference of the real part of eigenvalues $\lambda_{\nu}$ for $\nu\geq 1$ and $\lambda_{0}$ is always shown to be negative. The full spectrum can be checked for many different parameters in Fig. 5. This plot allows us to see the large dissipative gap and no oscillatory behavior in the NI phase (a3, a4, b4), as well as the nonzero imaginary components and a small gap in the $\mathcal{PT}$u phase (c1-3), and the intermediate gap and no oscillations in the $\mathcal{PT}$b phase (a1), in addition to the transitions between them.
@@
$b^{(0)}_{3}=b^{(0)}_{0}\left(1+\lambda_{0}\frac{\lambda_{0}-A}{2}\right),$ (32)
and $b^{(0)}_{1}=-b^{(0)}_{2}=-ib^{(0)}_{0}\lambda_{0}/2$. We choose $b^{(0)}_{0}$ such that the eigenvector is normalized for real $\lambda_{0}$, yielding the eigenvector associated to the stable steady state
@@
Appendix D Bloch sphere dynamics for the Stochastic Dissipative Qubit
@@
D.1 Polar coordinates: Nullclines
The Bloch sphere can be naturally parametrized in spherical coordinates $(r,\theta,\phi)$, related to the $(x,y,z)$ variables by the standard relations $x=r\sin\theta\cos\phi,\;y=r\sin\theta\sin\phi,\;z=r\cos\theta$. Using the chain rule, we can obtain a system of equations for the evolution of the polar variables as
$\begin{cases}\dot{r}=\Gamma\left((2\gamma\Gamma-1)(r^{2}-1)\cos\theta-\gamma\Gamma r\sin^{2}\theta\right),\\ \dot{\theta}=-\frac{\Gamma\sin\theta}{r}(1-2\gamma\Gamma+\gamma\Gamma r\cos\theta)-2J\sin\phi,\\ \dot{\phi}=-2J\cos\phi\cot\theta.\end{cases}$ (36)
Note that the similarity between the equation of motion for $r$ and for the purity, due to $P_{t}=(1+r_{t}^{2})/2$ yielding $\dot{P}_{t}=r_{t}\dot{r}_{t}$.
Let us now study the steady states of this system of equations. First, note that setting $\phi=\frac{\pi}{2}$, equivalent to $x=0$, trivially gives the steady state of the $\phi$ variable. In the study of nonlinear dynamics Strogatz (2018), the concept of nullclines, defined as the curves where $\dot{r}=0,\;\dot{\theta}=0,\;\dot{\phi}=0$, allows for a more intuitive understanding of the behavior of the dynamical system. This is because the steady states of the system live in the intersection of the nullclines. In particular, the nullcline associated to $r$, on which $\dot{r}=0$, describes the region where purity remains constant in Fig. 1 of the main text, given by
$r^{\textsc{n}_{r}}_{\theta}=\frac{\gamma\Gamma\sin^{2}\theta-\sqrt{4(1-2\gamma\Gamma)^{2}\cos^{2}\theta+\gamma^{2}\Gamma^{2}\sin^{4}\theta}}{(4\gamma\Gamma-2)\cos\theta}.$ (37)
This equation describes the white region of Fig. 1 of the main text with vanishing time-derivative of the purity, where the steady states live. Interestingly, it depends only on the dimensionless product $\gamma\Gamma$ but not on each of the variables independently. In the $\gamma\Gamma\rightarrow\infty$ limit this function converges to
$r^{\textsc{n}_{r}}_{\theta}=\frac{1-\cos(2\theta)-2\sqrt{16\cos^{2}\theta+\sin^{4}\theta}}{8\cos\theta}.$ (38)
The area enclosed by this curve is given by $\frac{1}{2}\int_{0}^{\pi}r_{\theta}^{2}\mathrm{d}\theta\approx 0.31\pi$, which means that, in this limit, $31\%$ of the states in the cut of the Bloch sphere are purifying.
In turn, the nullcline for $\theta$, i.e., the curve that lies in the $(y,z)$ plane of the Bloch sphere in which $\dot{\theta}=0$, is given by
$r^{\textsc{n}_{\theta}}_{\theta}=\frac{\Gamma(2\gamma\Gamma-1)\sin\theta}{2J+\gamma\Gamma^{2}\cos\theta\sin\theta}.$ (39)
This equation does depend on the dimensionless variables $\gamma J$ and $\Gamma/J$ separately and is used to obtain each of the points forming the green line in Fig. 1.
Appendix E Comparison of different numerical approaches
We now compare two different numerical resolutions of the SDQ dynamics. The first is to numerically integrate the nonlinear system of differential equations (35)—we do so by a standard 4th-order Runge-Kutta (RK4) method. The second approach is to get the time evolution of the density matrix from the formal solution of the master equation, i.e., $|\tilde{\rho}_{t})=e^{\tilde{\mathcal{L}}t}|\rho_{0})$, where the equation is vectorized to compute the map $e^{\tilde{\mathcal{L}}t}[\bullet]$ simply through matrix exponentiation. Additionally, this method requires normalization of the state $|\rho_{t})=|\tilde{\rho}_{t})/\mathrm{Tr}(\tilde{\rho}_{t})$, where $\mathrm{Tr}(\tilde{\rho}_{t})=(\mathbb{1}|\tilde{\rho}_{t})=\tilde{\rho}_{ee}+\tilde{\rho}_{ff}$ and $\tilde{\rho}_{ee}=\bra{e}\tilde{\rho}_{t}\ket{e}$. To avoid computing a matrix exponential for each time $t$, we Trotterize the evolution as $e^{\tilde{\mathcal{L}}t}=\prod_{n=1}^{N_{\textsc{t}}}e^{\tilde{\mathcal{L}}\Delta_{\textsc{t}}}$, subdividing the evolution in $N_{\textsc{t}}$ steps of length $\Delta_{\textsc{t}}=t_{\rm end}/N_{\textsc{t}}$.
These two approaches are compared in Fig. 7 to compute the purity, showing a perfect agreement. The first-order evolution of the purity obtained from $\partial_{t}P$ in eq. (9) of the main text, also agrees with the simulated evolution.
@@ Appendix E Comparison of different numerical approaches
Appendix F Uhlmann fidelity for a qubit
@@
This expression was used in Fig. 3 of the main text to compute the distinguishability between the instantaneous state and the stable steady state.
Appendix G Comparison of the anti-dephasing Liouvillian with hybrid and tilted Liouvillians
The Nonlinear Master Equation derived in the main text (4) is not of GKSL form and describes, to the best of our knowledge, an entirely new form of dissipation, what we call anti-dephasing. There are other master equations beyond GKSL form commonly used in the literature; we study their relations in this appendix.
The Hybrid Liouvillian Minganti et al. (2020) describes the dynamics of a system undergoing continuous monitoring of quantum jumps and post-selection, with a finite-efficiency $\eta$ detector, and reads, for a single jump operator $\hat{L}$ with rate $\mu$
~$\tilde{\mathcal{L}}_{q}[\bullet]=-i[\hat{H}_{0},\bullet]+\mu(q\hat{L}\bullet\hat{L}^{\dagger}-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\bullet\}),$ (42)
which nicely interpolates between Non-Hermitian evolution when the detector is totally efficient $q=1-\eta=0$ and Lindblad dynamics when the detector is totally inefficient $q=1-\eta=1$. This hybrid Liouvillian only has a physical interpretation in terms of post-selected trajectories when $q\in[0,1]$. The dynamics is always trace-decreasing when $q\in[0,1)$ and trace-preserving for $q=1$.
Another commonly considered master equation beyond GKSL form is the tilted generator Garrahan and Lesanovsky (2010); Carollo et al. (2018) or generalized quantum master equation Esposito et al. (2009), whose classical analog is the Lebowitz-Spohn operator Lebowitz and Spohn (1999). This generator describes the dynamics of a biased ensemble of trajectories and, in its simplest form, reads
~$\tilde{\mathcal{L}}_{s}[\bullet]=-i[\hat{H},\bullet]+\mu(e^{-s}\hat{L}\bullet\hat{L}^{\dagger}-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\bullet\}).$ (43)
This generator describes the dynamics of the biased ensemble of trajectories $\tilde{\rho}_{s}(t)=\sum_{K=0}^{\infty}\tilde{\rho}^{(K)}(t)e^{-sK}$, where $\tilde{\rho}^{(K)}$ represents the density matrix of the dynamics with $K$ events after time $t$, i.e. jumps with operator $\hat{L}$. The variable $s$ represents the conjugate field to $K$. The dynamics generated by this equation is not trace-preserving if $s\neq 0$, if $s<0$ the system is in the active phase, in which the trajectories with jumps are favored, and the dynamics is trace-increasing, and if $s>0$ the system is in the passive phase with less jumps than usual, and the dynamics is trace-decreasing.
Both of these generators do not alter the form of the anticommutator term in the GKSL master equation $-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\bullet\}$. The nonlinear master equation derived in this work, equation (4) of the main text, contains a double anticommutator, which means that this term is positive, instead of negative, as in GKSL. For this reason, the nonlinear master equation does not simply reduce to one of the previously mentioned generators.
There is a particular case in which the connection is closer, if the jump operator is proportional to a projector, i.e., $\hat{L}\propto\hat{\Pi}$, where $\hat{\Pi}^{2}=\hat{\Pi}$, as in the SDQ example studied in the main text (8). In this case, through the mapping $\Gamma-\gamma\Gamma^{2}\equiv\tfrac{1}{2}\mu$ and $2\gamma\Gamma^{2}\equiv\mu e^{-s}$ the nonlinear master equation for the qubit can be interpreted as a tilted generator and with $2\gamma\Gamma^{2}\equiv\mu q$ it can be interpreted as a hybrid Liouvillian. However, the two mappings are only valid when the anticommutator term is negative $\Gamma-\gamma\Gamma^{2}>0$, i.e., $\gamma<\Gamma^{-1}$. Furthermore, the mapping to a hybrid Liouvillian has no physical interpretation when the dynamics is trace increasing, i.e. $\gamma>\frac{1}{2\Gamma}$.
Appendix H Standard form of the anti-dephasing master equation
@@
where $\tilde{\mathcal{L}}$ need not be trace-preserving. The normalized density matrix $\hat{\rho}$ evolves according to
@@
However, by relaxing this condition, the general structure of ~$\tilde{\mathcal{L}}[\hat{\rho}]$ is considered. The rate of norm change, in this case, is
@@
$\frac{d}{dt}\hat{\rho}-i[\hat{H},\hat{\rho}]+\{\hat{G},\hat{\rho}\}+\sum_{ij=1}^{N^{2}}a_{ij}\hat{F}_{i}\hat{\rho}\hat{F}_{j}^{\dagger}-\Big{(}2{\rm Tr}(\hat{G}\hat{\rho})+\sum_{ij=1}^{N^{2}}a_{ij}{\rm Tr}(\hat{F}_{j}^{\dagger}\hat{F}_{i}\hat{\rho})\Big{)}\hat{\rho}.$ (51)
@@
$\frac{d}{dt}\hat{\rho}=-i[\hat{H},\hat{\rho}]+\{\hat{G},\hat{\rho}\}+\sum_{ij=1}^{N^{2}}\gamma_{k}\hat{A}_{k}\hat{\rho}\hat{A}_{k}^{\dagger}-\Big{(}2{\rm Tr}[\hat{G}\hat{\rho}]+\sum_{k=1}^{N^{2}}\gamma_{k}{\rm Tr}[\hat{A}_{k}^{\dagger}\hat{A}_{k}\hat{\rho}]\Big{)}\hat{\rho},$ (52)